OpenStudy (anonymous):

Could somebody PLEASE help me find 3 pythagorean theorem proofs and give a short explanation for each one? Please!!!

OpenStudy (anonymous):

The pythagorean theorem is a^2 + b^2 = c^2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so... 3, 4, 5 work for the pythagorean theorem... 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25

OpenStudy (anonymous):

http://lmgtfy.com/?q=pythagoras+proof

OpenStudy (anonymous):

5, 12, 13 8, 15, 17

OpenStudy (anonymous):

think you are looking for a proof of the theorem there are many some take algebra, some just pictures

OpenStudy (anonymous):

@satellite73, nice :D